Given that ˝ and -3 are the roots of the equation ax˛+bx+c=0. Find a,b and c.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Given that ˝ and -3 are the roots of the equation ax˛+bx+c=0. Find a,b and c.
--------
(x - 1/2)*(x + 3) = 0
(2x - 1)*(x + 3) = 0
2x^2 + 5x - 3 = 0
--------
a = 2, b = 5, c = -3
-------------------
a, b & c can be any multiple of the coefficients listed.
(2,5,-3)*k

Answer by stanbon(75887)

Given that ˝ and -3 are the roots of the equation ax˛+bx+c=0. Find a,b and c.
---
sum = -b/a = -5/2
product = c/a = -3/2
---------
Looks like a = 2, b = 5 and c = -3
------------------
You can find these formulas by adding and by multiplying
the roots shown in the Quadratic Formula.
----------------------
Cheers,
Stan H.
------------